OpenStudy (anonymous):

using the formula y=mx=b, find the equation of the line determined by the two points (3,-8) and (9,5)

OpenStudy (anonymous):

First you have to find the slope between the two points

OpenStudy (anonymous):

Do you know how to find slope?

OpenStudy (anonymous):

(y2-y1)/(x2-x1)

OpenStudy (anonymous):

yes but i think it wants us to plug in the -8 in for y and 3 in for x, the the next set for the next equation

OpenStudy (anonymous):

not yet, first you have to find slope

OpenStudy (anonymous):

slope is 1/2?

OpenStudy (anonymous):

y=1/2x+b now you plug in your y and x value to find b

OpenStudy (anonymous):

of either point

OpenStudy (anonymous):

would it be -8?

OpenStudy (anonymous):

-8=1/2(3)+b -8 - 3/2=b

OpenStudy (anonymous):

the answer was 13/6x+-29/2